OpenStudy (anonymous):

WILL MEDAL Enter The Answer: What is \[\frac{ (n-5)*6 }{ n÷4 }\]When n=8

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Time to take out your calculator if you cannot figure out without one, but do remember the rules of PEMDAS.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

its 9

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

bro do it like this 8-5=3 3x6=18 8/4=2 18/2=9

OpenStudy (anonymous):

thanks i fanned and medaled u

OpenStudy (anonymous):

np, i did it back